# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

### v7

@@ 3.2 Quadratic Gradient Definition
Suppose that a differentiable scalar-valued function $F(\mathbf{x})$ has its gradient $\boldsymbol{g}$ and Hessian matrix $H$, with any matrix $\bar{H}\leq H$ in the Loewner ordering for a maximization problem as follows:
$\boldsymbol{g}=\left[\begin{array}[]{c}g_{0}\\ g_{1}\\ \vdots\\ g_{d}\\ \end{array}\right],$ $H=\left[\begin{array}[]{cccc}\nabla_{00}^{2}&\nabla_{01}^{2}&\ldots&\nabla_{0d}^{2}\\ \nabla_{10}^{2}&\nabla_{11}^{2}&\ldots&\nabla_{1d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \nabla_{d0}^{2}&\nabla_{d1}^{2}&\ldots&\nabla_{dd}^{2}\\ \end{array}\right],\quad$
$\bar{H}$ $=\left[\begin{array}[]{cccc}\bar{h}_{00}&\bar{h}_{01}&\ldots&\bar{h}_{0d}\\ \bar{h}_{10}&\bar{h}_{11}&\ldots&\bar{h}_{1d}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{h}_{d0}&\bar{h}_{d1}&\ldots&\bar{h}_{dd}\\ \end{array}\right],$
@@
Given such a $\tilde{B}$ above, we define the quadratic gradient as $G=\bar{B}\cdot\boldsymbol{g}$ with a new learning rate $\eta$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$, and $\eta$ should be always no less than 1 and decrease to 1 in a limited number of iteration steps. Note that $G$ is still a column vector of the same size as the gradient $\boldsymbol{g}$. To maximize the function $F(\mathbf{x})$, we can use the iterative formulas: $\mathbf{x}_{t+1}=\mathbf{x}_{t}+\eta\cdot G$, just like the naive gradient.
To minimize the function $F(x)$ is the same as to just maximize the function $-F(x)$, in which case we need to construct the $\tilde{B}$ by any good lower bound $\bar{H}$ of the Hessian $-H$ of $-F(x)$ or any good upper bound $\bar{H}$ of the Hessian $H$ of $F(x)$. We point out here that $\bar{H}$ could be the Hessian matrix $H$ itself.
@@
1:  Require: training dataset $X\in\mathbb{R}^{n\times(1+d)}$; training label $Y\in\mathbb{R}^{n\times 1}$; learning rate $lr\in\mathbb{R}$; and the number $\kappa$ of iterations;
2:  Ensure: the parameter vector $V\in\mathbb{R}^{(1+d)}$
3:  Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$ $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
4:  Set $V\leftarrow\boldsymbol{0}$, $W\leftarrow\boldsymbol{0}$, $\bar{B}\leftarrow\boldsymbol{0}$ $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
5:  for $i:=0$ to $d$ do
6:     # $\epsilon$ is a small positive constant such as $81e-$ 8
7:     $\bar{B}[i][i]\leftarrow\epsilon$
8:     for $j:=0$ to $d$ do
9:        $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
10:     end for
11:  end for
12:  Set $alpha_{0}\leftarrow 0.01$
13:  Set $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
14:  for $count:=1$ to $\kappa$ do
15:     # $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
16:     Set $Z\leftarrow\boldsymbol{0}$
17:     for $i:=1$ to $n$ do
18:        for $j:=0$ to $d$ do
19:           $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
20:        end for
@@
22:     # $\boldsymbol{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
23:     Set $\boldsymbol{\sigma}\leftarrow\boldsymbol{0}$
@@
25:        $\boldsymbol{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
@@
27:     Set $\boldsymbol{g}\leftarrow\boldsymbol{0}$
28:     for $j:=0$ to $d$ do
29:        for $i:=1$ to $n$ do
30:           $\boldsymbol{g}[j]\leftarrow\boldsymbol{g}[j]+(1-\boldsymbol{\sigma}[i])\times Y[i]\times X[i][j]$
@@
32:     end for
33:     Set $G\leftarrow\boldsymbol{0}$
34:     for $j:=0$ to $d$ do
35:        $G[j]\leftarrow\bar{B}[j][j]\times\boldsymbol{g}[j]$
36:     end for
37:     Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$
38:     Set $\gamma\leftarrow lr/(n\times count)$
39:     for $j:=0$ to $d$ do
40:        $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$
41:        $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$
42:        $W[j]\leftarrow w_{temp}$
43:     end for
44:     $alpha_{0}\leftarrow alpha_{1}$
45:     $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
46:  end for
47:  Return $V$
@@ 3.3 Quadratic Gradient Algorithms
This gradient variant, Quadratic Gradient, can be used to enhance various first-order gradient algorithms:
@@
    $V_{t+1}$ $=\boldsymbol{\beta}_{t}+\eta_{t}\cdot\nabla J(\boldsymbol{\beta}_{t}),$ (3)
    $\boldsymbol{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\boldsymbol{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position (Kim et al., 2018a). The enhanced NAG is to replace (3) with $V_{t+1}=\boldsymbol{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced algorithm usually setted to $1+\eta_{t}$. Our enhanced NAG is described in Algorithm 1 .
@@
    Adagrad is a gradient-based algorithm suitable for dealing with sparse data. The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\boldsymbol{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}\boldsymbol{g}_{[i]}^{(t)}\cdot\boldsymbol{g}_{[i]}^{(t)}}}\cdot\boldsymbol{g}_{[i]}^{(t)},$
    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
@@
    Adam is an optimization algorithm that combines the benefits of Adagrad, which adapts learning rates for each parameter, with momentum, allowing it to converge faster and handle sparse gradients more effectively. Similar to how Adagrad transitions to its enhanced versions, the update formula for enhanced Adam, also employs a new learning rate and replaces the gradient with its quadratic form.
@@
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 1: The training results of NAG and Enhanced NAG in the clear
@@
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 2: The training results of Adagrad and Enhanced Adagrad in the clear
@@
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 3: The training results of Adam and Enhanced Adam in the clear
@@
Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM. Given that our focus is on the convergence speed of the algorithms during training, we use the loss function, specifically maximum likelihood estimation (MLE), as the only indicator. We evaluate six algorithms—NAG, Adagrad, Adam, and their quadratic-gradient variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam, respectively)—using the datasets adopted by Kim et al. (2018a): the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (Edinburgh), the Low Birth Weight Study (lbw), Nhanes III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study datasets (uis). The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer. The remaining five datasets each contain a single binary dependent variable. We also evaluate these algorithms on two large datasets from Han et al. (2019): a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features. For a fair comparison with the baseline work (Kim et al., 2018a), we use the learning rate $1+\frac{10}{1+t}$, in line with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate. In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the binary logistic regression model.
Figures 21, and 3 show that, except for the Enhanced Adagrad and Adam methods on the iDASH genomic dataset, our enhanced methods converge faster than their original counterparts in all other cases. All figures demonstrate that the enhanced algorithms converge empirically. In all the Python experiments, the time required to calculate $\bar{B}$ in the quadratic gradient $G$ before running the iterations, as well as the time to execute each iteration for various algorithms, is minimal and can be considered negligible (a few seconds).
@@
An important observation is that the enhanced algorithms have a better performance when using learning rates between $1$ and $2$ than using learning rate otherelse. When using a learning rate greater than $3$, the algorithm is almost certain not to converge. When employing quadratic gradient algorithms, we recommend adopting the learning rate setting $1+A\cdot\gamma^{t}$ as the optimal choice, where $t$ represents the iteration number, $A$ is a positive float number typically set to no less than $1$, and $\gamma$ is usually a positive number less than 1, controlling the rate of decay. We strongly recommend that $A$ is set to $1.0$ and $\gamma$ is tuned to search for optimal results, or that one can set $A$ to a small number less than $1.0$ and $\gamma$ to a larger number greater than $0.9$.
Analysis In Figure 2(a), the enhanced Adagrad algorithm failed to outperform the original Adagrad algorithm. The possible reason for that might be related to the limitations of the raw Adagrad method. Without a doubt, Adagrad is a novel algorithm initiated to accelerate each element of the gradient with different learning rates. However, Adagrad tends to converge to a suboptimal solution due to its aggressive, monotonically decreasing learning rates. This would lead to its main limitation that in the later training phase every learning rate for different components of the gradient is too close to zero due to keeping adding positive additional terms to the denominator, stopping the algorithm from learning anything.
On the other hand, the original Adagrad method has another little-noticed limitation: the learning rate in the first few iterations tends to be large. While this limitation does not affect the performance of the original Adagrad method to some extent, the enhanced Adagrad method exacerbates this phenomenon significantly, leading to the Learning-Rate Explosion . Therefore, the enhanced Adagrad cannot be applied to general optimization problems such as Rosenbrock’s function. The exploding learning rate would be too large for the algorithm to survive the first several iterations, finally leading the optimization function to some point where its output cannot be represented by the computer system. This might explain why the performance of this algorithm in all cases, not just on the iDASH genome dataset, seems to be meaningless and unstable in the first few iterations.
Several improved algorithms upon the Adagrad method, such as RMSProp, have been proposed in order to address these issues existed, via using an exponential moving average of historical gradients rather than just the sum of all squared gradients from the beginning of training. We might be able to overcome the problems existing in the enhanced Adagrad method by adopting the enhanced Adagrad-like variants, like the enhanced Adadelta method and the enhanced RMSProp method.
The reasons why the enhanced algorithms have a better performance might be that quadratic gradient can incorporate curves into first-order gradient algorithms.
(a) The raw first-order gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{lr}\cdot\boldsymbol{g}$
(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{LR}\cdot G$
Figure 4: The training outcomes for both the raw gradient algorithm and the raw quadratic gradient ascent algorithm in the clear setting, conducted on the lbw dataset
@@
Gradient And Quadratic Gradient We executed the raw gradient ascent algorithm and raw quadratic gradient algorithm on the lbw dataset using various learning rates. Figure 4 shows the detailed results of this experiment. It is precisely these results that inspired the present authors to come up with the idea of quadratic gradient. Similar to the gradient, the quadratic gradient carrying the curve also exhibits the characteristic of smooth progression. That is, when the learning rate is gradually increased, both the gradient and the quadratic gradient will show corresponding gradual changes in performance rather than abrupt shifts.
4 Homomorphic Training

### v8

@@ 3.2 Quadratic Gradient Definition
Suppose that a differentiable scalar-valued function $F(\mathbf{x})$ has its gradient $\bm{g}$ and Hessian matrix $H$, with any matrix $\bar{H}\leq H$ in the Loewner ordering for a maximization problem as follows:
$\bm{g}=\left[\begin{array}[]{c}g_{0}\\ g_{1}\\ \vdots\\ g_{d}\\ \end{array}\right],$ $H=\left[\begin{array}[]{cccc}\nabla_{00}^{2}&\nabla_{01}^{2}&\ldots&\nabla_{0d}^{2}\\ \nabla_{10}^{2}&\nabla_{11}^{2}&\ldots&\nabla_{1d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \nabla_{d0}^{2}&\nabla_{d1}^{2}&\ldots&\nabla_{dd}^{2}\\ \end{array}\right],\quad\bar{H}$ $=\left[\begin{array}[]{cccc}\bar{h}_{00}&\bar{h}_{01}&\ldots&\bar{h}_{0d}\\ \bar{h}_{10}&\bar{h}_{11}&\ldots&\bar{h}_{1d}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{h}_{d0}&\bar{h}_{d1}&\ldots&\bar{h}_{dd}\\ \end{array}\right],$
@@
Given such a $\tilde{B}$ above, we define the quadratic gradient as $G=\bar{B}\cdot\bm{g}$ with a new learning rate $\eta$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$, and $\eta$ should be always no less than 1 and decrease to 1 in a limited number of iteration steps. Note that $G$ is still a column vector of the same size as the gradient $\bm{g}$. To maximize the function $F(\mathbf{x})$, we can use the iterative formulas: $\mathbf{x}_{t+1}=\mathbf{x}_{t}+\eta\cdot G$, just like the naive gradient.
Minimizing the function $F(x)$ is equivalent to maximizing the function $-F(x)$. In this context, we need to construct $\tilde{B}$ using either a good $lower$ bound $\bar{H}$ for the Hessian $-H$ of $-F(x)$ or a good $upper$ bound $\bar{H}$ for the Hessian $H$ of $F(x)$. It is worth noting that $\bar{H}$ can be taken as the Hessian $H$ itself, thereby eliminating the need to find either lower or upper bounds.
Böhning and Lindsay Böhning & Lindsay (1988) did not proposed a systemic way to find or build the constant Hessian approximate for their fixed Hessian method. Perhaps there are no such matrices for most objective functions. We mention here that our quadratic gradient doesn’t need a constant (fixed) Hessian replacement to construct and that we can use the original Hessian itself to build our varying diaginal matrix $\tilde{B}$ for the quadratic gradient. Because of this loose condition, we propose a system way satisfying sufficient condition to find if there is a fixed Hessian matrix for efficient calculation. That is, we first build $\tilde{B}$ from the Hessian of the given objective function and then check if for every its diaginal element $\tilde{B}_{kk}$ there is a constant maximum for $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$ (or constant minimum $|\tilde{B}_{kk}|$). If each diagnal element $|\tilde{B}_{kk}|$ all had their minimum $|\tilde{M}_{kk}|$, then the constant diagnial matrix with corresponding diagnal elements $|\tilde{M}_{kk}|$ would satisfy the convengence condiion for SFH method. Finally, we should also evaluate the performance of such fixed Hessian replacement since it might not be a $good$ bound. This searching method can be employed to any optimization functions. For example, we can build $|\tilde{B}_{kk}|$ from the Hessian $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$ for the logistic regressian and attempt to maximise $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$. The resulting constant diagnoal matrix is the same as the one built from the $-\frac{1}{4}X^{\top}X$ using the quadratic-gradient construction.
Böhning and Lindsay Böhning & Lindsay (1988) did not propose a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method. This may be due to the absence of such matrices for most objective functions. We note that our quadratic gradient approach does not require a constant (fixed) Hessian replacement and instead allows for the use of the original Hessian itself to construct our varying diagonal matrix $\tilde{B}$ for the quadratic gradient. Under this less restrictive condition, we present a methodical approach to evaluate whether a fixed Hessian matrix can be employed for efficient computation. Specifically, we first construct $\tilde{B}$ from the Hessian of the given objective function and then check if, for each diagonal element $\tilde{B}_{kk}$, there is a constant maximum for $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$ (or a constant minimum $|\tilde{B}_{kk}|$). If each diagonal element $|\tilde{B}_{kk}|$ achieves its minimum $|\tilde{M}_{kk}|$, then a constant diagonal matrix with the corresponding diagonal elements $|\tilde{M}_{kk}|$ would satisfy the convergence condition for the SFH method. Finally, we should evaluate the performance of such a fixed Hessian replacement, as it might not always provide a good bound. This search method can be applied to any optimization function to identify a fixed Hessian. However, the absence of a constant Hessian replacement via this method does not necessarily imply that such an approximation does not exist.
Our method can be employed to identify a fixed Hessian for binary logistic regression. For instance, we can construct $|\tilde{B}_{kk}|$ from the Hessian $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$ and then try to maximize $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$. The resulting constant diagonal matrix is equivalent to the one obtained from $-\frac{1}{4}X^{\top}X$ using the quadratic-gradient construction. Therefore, $-\frac{1}{4}X^{\top}X$ could serve as the fixed Hessian for the fixed Hessian method, precisely as suggested by Böhning and Lindsay Böhning & Lindsay (1988).
@@
1:training dataset $X\in\mathbb{R}^{n\times(1+d)}$; training label $Y\in\mathbb{R}^{n\times 1}$; learning rate $lr\in\mathbb{R}$; and the number $\kappa$ of iterations;
2:the parameter vector $V\in\mathbb{R}^{(1+d)}$
3:Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$, $V\leftarrow\bm{0}$, $W\leftarrow\bm{0}$, $\bar{B}\leftarrow\bm{0}$
4:for $i:=0$ to $d$ do $\triangleright$ $\epsilon$ is a small positive constant such as $81e-$ 8
5:     $\bar{B}[i][i]\leftarrow\epsilon$
6:     for $j:=0$ to $d$ do
7:         $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
8:     end for
9:end for
10:Set $alpha_{0}\leftarrow 0.01$, $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
11:for $count:=1$ to $\kappa$ do
12:     Set $Z\leftarrow\bm{0}$triangleright $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
13:     for $i:=1$ to $n$ do
14:         for $j:=0$ to $d$ do
15:              $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
16:         end for
17:     end for
18:     Set $\bm{\sigma}\leftarrow\bm{0}$triangleright $\bm{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
19:     for $i:=1$ to $n$ do
20:         $\bm{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
@@
22:     Set $\bm{g}\leftarrow\bm{0}$
23:     for $j:=0$ to $d$ do
@@
25:              $\bm{g}[j]\leftarrow\bm{g}[j]+(1-\bm{\sigma}[i])\times Y[i]\times X[i][j]$
@@
27:     end for
28:     Set $G\leftarrow\bm{0}$
29:     for $j:=0$ to $d$ do
30:         $G[j]\leftarrow\bar{B}[j][j]\times\bm{g}[j]$
@@
32:     Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$, $\gamma\leftarrow(\frac{10}{1+t})/n$
33:     for $j:=0$ to $d$ do
34:         $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$, $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$, $W[j]\leftarrow w_{temp}$
35:     end for
36:     $alpha_{0}\leftarrow alpha_{1}$, $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
37:end for
38:Return $V$
@@ 3.3 Quadratic Gradient Algorithms
Quadratic gradient can be used to enhance various first-order gradient algorithms:
@@
    $V_{t+1}$ $=\bm{\beta}_{t}+\eta_{t}\cdot\nabla J(\bm{\beta}_{t}),$ (3)
    $\bm{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\bm{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position (Kim et al., 2018a). The enhanced NAG is to replace equation 3 with $V_{t+1}=\bm{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced algorithm usually setted to $1+\eta_{t}$. Our enhanced NAG is described in Algorithm 1 .
@@
    Adagrad is a gradient-based algorithm suitable for dealing with sparse data. The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\bm{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
    $\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}\bm{g}_{[i]}^{(t)}\cdot\bm{g}_{[i]}^{(t)}}}\cdot\bm{g}_{[i]}^{(t)},$
    $\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
@@
    Adam is an optimization algorithm that combines the benefits of Adagrad, which adapts learning rates for each parameter, with momentum, allowing it to converge faster and handle sparse gradients more effectively. Similar to how Adagrad transitions to its enhanced version, the update formula for enhanced Adam, also employs a new learning rate $N_{t}$ and replaces the gradient with its quadratic form.
@@
(g) The MNIST dataset
(h) The financial dataset
Figure 1: The training results of NAG and Enhanced NAG in the clear.
@@
(g) The MNIST dataset
(h) The financial dataset
Figure 2: The training results of Adagrad and Enhanced Adagrad in the clear.
@@
(g) The MNIST dataset
(h) The financial dataset
Figure 3: The training results of Adam and Enhanced Adam in the clear.
@@
Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM. Given that our focus is on the convergence speed of the algorithms during training, we use the loss function $l(\bm{\beta})$, specifically maximum likelihood estimation (MLE), as the only indicator. We evaluate six algorithms—NAG, Adagrad, Adam, and their quadratic-gradient variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam, respectively)—using the datasets adopted by Kim et al. Kim et al. (2018a): the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (edin), the Low Birth Weight Study (lbw), Nhanes III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study datasets (uis). The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer. The remaining five datasets each contain a single binary dependent variable. We also evaluate these algorithms on two large datasets from Han et al. Han et al. (2019): a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features. For a fair comparison with the baseline work (Kim et al., 2018a), we use the learning rate $1+\frac{10}{1+t}$, in line with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate. The enhanced Adagrad algorithm use $N_{t}=1+0.01$ as the learning rate while the enhanced Adam algorithms adopt the following learning rate setting: $\alpha=1+0.001,$ $\beta_{1}=0.9,$ and $\beta_{2}=0.999.$ In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\bm{\beta})$.
From the empirical results presented in Figures 12, and 3, we conclude that all the enhanced algorithms converge, and that the enhanced NAG algorithms outperform the original NAG method in terms of convergence rate. Although the enhanced Adagrad and Adam methods do not achieve faster convergence than their original counterparts on the iDASH genomic dataset, they demonstrate clear advantages in all other cases. The reason for this discrepancy is beyond the scope of this paper and will be addressed as part of future work. One possible explanation for the superior performance of the enhanced algorithms is that the quadratic gradient integrates curvature information into first-order gradient methods.
@@
An important observation is that the enhanced algorithms demonstrate better performance with learning rates between $1$ and $2$ compared to other values. When the learning rate exceeds $3$, the algorithm is nearly guaranteed not to converge. For quadratic gradient algorithms, we recommend employing an exponentially decaying learning rate, such as $1+A\cdot\gamma^{t}$, where $t$ denotes the iteration number, $A$ is a positive constant typically set to no less than $1$, and $\gamma$ is a positive number less than $1$ that controls the rate of decay.
@@
Gradient And Quadratic Gradient We executed the raw gradient ascent algorithm and the raw quadratic gradient algorithm on the LBW dataset using various learning rates. Figure 4 presents the detailed results of this experiment. It was precisely these results that directly inspired the authors to develop the concept of the quadratic gradient. Similar to the gradient, the quadratic gradient, which incorporates curvature, also demonstrates smooth progression. In particular, as the learning rate is gradually increased, both the gradient and the quadratic gradient reveal corresponding gradual changes in performance, rather than sudden jumps.
(a) The raw first-order gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{lr}\cdot\bm{g}$
(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{LR}\cdot G$
Figure 4: The training outcomes for both the raw gradient algorithm and the raw quadratic gradient ascent algorithm in the clear setting, conducted on the lbw dataset.
4 Secure Training